OpenStudy (gabylovesyou):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 19 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (gabylovesyou):

@hba

OpenStudy (hba):

Which method do you prefer ?

OpenStudy (gabylovesyou):

which one is easier

OpenStudy (hba):

If i would be in your place i would be using elimination.

OpenStudy (gabylovesyou):

ok then elimination

OpenStudy (hba):

So we have two equation in this question :) x - y = 19---(1) 2x + y = 2---(2) 1)Multiply equation (1) by 2 :)

OpenStudy (gabylovesyou):

x - y = 38

OpenStudy (hba):

2(x-y)=2(19)

OpenStudy (gabylovesyou):

what i did was wrong? or you want me to do that ^

OpenStudy (gabylovesyou):

2(x - y) = 38

OpenStudy (hba):

Yes we have to multiply 2 by both sides.

OpenStudy (hba):

So, 2x-2y=38---(3) Now subtract it from Eq(2) =)2x + y = 2 2x - 2y=38 - + - --------------

OpenStudy (gabylovesyou):

i subtract 2 - 38??

OpenStudy (hba):

Yes both the sides.

OpenStudy (hba):

If you are not getting it let us use the substitution method.

OpenStudy (hba):

You can write in the second part:I don't know elimination.

OpenStudy (gabylovesyou):

LOL nooo i think the answer is x = 7, y = - 12 IF I GET IT RIGHT THEN I DO lol

OpenStudy (hba):

Yeah right.

OpenStudy (gabylovesyou):

im right??

OpenStudy (hba):

Yes.

OpenStudy (gabylovesyou):

ok can you put everything i did together so i put it in part 1

OpenStudy (hba):

I cannot do the whole solution for you.

OpenStudy (hba):

Btw,If you have done it on your own,You may know the solution ??

OpenStudy (gabylovesyou):

i just told u the solution lol can you just write the steps??

OpenStudy (hba):

No.

OpenStudy (hba):

Stop playing gaby.

OpenStudy (gabylovesyou):

whyyyyyyyyyyyy its my last question -.- pleaseeeeeee

OpenStudy (gabylovesyou):

please

OpenStudy (gabylovesyou):

please

OpenStudy (gabylovesyou):

please

OpenStudy (hba):

Ahm,Ahm....Ahm

OpenStudy (gabylovesyou):

last question and i dont bother u no more lol please tell me the steps... i already have the solution but i need to show my work

OpenStudy (hba):

So do as I say... I cannot do all the working. Sorry :/

OpenStudy (gabylovesyou):

oh my LORDDDDDDD just check this: is this good for part 1?x-y=19 2x+y=2 --------- 3x=21 (Add the two equations term by term.) x=7 (divide by 3) Now, solve for y using substitution. x-y=19 7-y=19 -y=12 (Subtract 7 from both sides.) y=-12

OpenStudy (hba):

Where did you get 3x=21 from ?

OpenStudy (hba):

The rest is fine :)

OpenStudy (gabylovesyou):

what do i do to fix the steps?

OpenStudy (gabylovesyou):

@hba DONT LEAVE MEEEE lol

OpenStudy (hba):

I am out of cookies :/

OpenStudy (gabylovesyou):

Part 1: x-y=19 2x+y=2 ____________ x=7 x-y=19 7-y=19 -y=12 y=-12 Part 2: I chose Elimination because I feel like its easier.

OpenStudy (hba):

C is for cookie and cookie is for me.

OpenStudy (gabylovesyou):

Is that a good answer? ^^^

OpenStudy (hba):

Lemme see.

OpenStudy (hba):

Show this before x=7 3x=21

OpenStudy (gabylovesyou):

WHAT DIDNT U SAY .. -.- ok lol

OpenStudy (hba):

kay.

OpenStudy (hba):

Do i get a cookie for this ?